# A non-parametric approach for estimating consumer valuation distributions using second price auctions

### v2

@@ 2.2 Likelihood based on the observed data from multiple identical, non-overlapping second price auctions
Suppose we consider $K$ independent second price auctions of identical copies of an item (with possibly different reserve prices $r_{1},r_{2},\ldots,r_{K}$). The observed data is $\{(M_{k},O_{k})\ ,\ \{(X_{i,k}\ ,\ T_{i-1,k})\}_{i=1}^{M_{k}}\}_{k=1}^{K}$, where $M_{k}$ denotes the number of selling price changes for the $k^{th}$ auction, $O_{k}$ denotes the indicator of the item being sold at the end of the $k^{th}$ auction, $\{X_{i,k}\}_{i=1}^{M_{k}}$ denote the selling/standing price sequence for the $k^{th}$ auction, and $\{T_{i-1,k}\}_{i=1}^{M_{k}}$ denote the sequence of intermediate waiting times between successive changes in the standing prices. Finally, let $T_{M_{k},k}=\tau-\sum_{i=0}^{M_{k}-1}T_{i,k}$, for all $k=1,2,\ldots,K$. Since the auctions are independent, it follows by Lemma 2.1 that the likelihood function of the unknown parameters $\lambda$ and $F$ for observed data values $\{(m_{k},o_{k})\ ,\ \{(x_{i,k}\ ,\ t_{i-1,k})\}_{i=1}^{m_{k}}\}_{k=1}^{K}$ is given by $Lik(\lambda,F)=\prod_{k=1}^{K}g(m_{k},o_{k},\{x_{i,k}\}_{i=1}^{m_{k}},\{t_{i,k}\}_{i=0}^{m_{k}-1})$.
@@
Note that the function $F$ is constrained to be non-decreasing. A key transformation to a constraint-free parametrization (described in Section 2.4 below) is needed to facilitate the conditional maximization of $Lik$ with respect to $F$. A crucial precursor to this re-parametrization is introduction of some new notation obtained by merging the standing prices from all the $K$ different auctions together. Let $L=\sum_{k=1}^{K}M_{k}$ denote the total number of standing/selling price changes in all the $K$ auctions. Recall that $\{(m_{k},o_{k})\ ,\ \{(x_{i,k}\ ,\ t_{i-1,k})\}_{i=1}^{m_{k}}\}_{k=1}^{K}$ denote the observed data values, and $t_{m_{k},k}=\tau-\sum_{i=0}^{m_{k}-1}t_{i,k}$ for $1\leq k\leq K$. Let $\ell$ denote the observed value of $L$. We will denote by ${\bf\bar{x}}=(\bar{x}_{1},\bar{x}_{2},\cdots,\bar{x}_{\ell})$ the arrangement/ordering of the pooled collection $\{\{x_{i,k}\}_{i=1}^{m_{k}}\}_{k=1}^{K}$ such that $\bar{x}_{1}<\bar{x}_{2}<\ldots<\bar{x}_{\ell}$; under the assumption that $F$ is a continuous cdf, there should be no ties in the entries of $\bf\bar{x}$} with probability one. In other words, we pool the standing prices from all the auctions (excluding the reserve prices) and then arrange them in ascending order as $(\bar{x}_{1},\bar{x}_{2},\cdots,\bar{x}_{\ell})$. Also, for $1\leq i\leq L$, we define $\bar{t}_{i}=t_{\bar{i},\bar{k}}$ where $\bar{i}$ and $\bar{k}$ are such that $\bar{x}_{i}=x_{\bar{i},\bar{k}}$. Further, let
@@
item is sold above the reserve price,
@@
Example (continued): Consider the example in Figure 2, where we have a setting with $K=4$ independent second price auctions, with reserve prices $(r_{1},r_{2},r_{3},r_{4})=(10,5,13,17)$. The first auction has $m_{1}=3$ standing price changes with $(x_{1,1},\ x_{2,1},\ x_{3,1})=(12,15,19)$, the second auction has $m_{2}=4$ standing price changes with $(x_{1,2},\ x_{2,2},\ x_{3,2},\ x_{4,2})=(16,18,20,25)$, the item is sold at the reserve price $r_{3}=13$ in the third auction ($m_{3}=0,o_{3}=1$), and the item is unsold in the fourth auction ($m_{4}=0,o_{4}=1$). Pooling and rearranging the standing prices (excluding reserve prices) from all the auctions, we see that $\ell=3+4+0+0=7$, and
@@
Note that the auction item is sold above the reserve price in the first two auctions, and the final selling prices are $x_{m_{1},1}=19$ and $x_{m_{2},2}=25$ respectively. Examining the positions of these two prices in $\bf\bar{x}$} gives us $S=\{5,7\}$. Finally, $K_{s}=\{1,2,3\}$ is the collection of auction indices where the item is sold.
@@
where $\bar{x}_{0}=0$. We now establish a final bit of notation necessary to introduce the constraint-free reparametrization of $F$. We now pool the $\ell+K$ standing prices from all the auctions (including the reserve prices), i.e., $\{\{x_{i,k}\}_{i=0}^{m_{k}}\}_{k=1}^{K}$, and arrange them in ascending order as $z_{1}<z_{2}<\cdots<z_{\ell+K}$, and denote ${\bf z}:=(z_{1},z_{2},\cdots,z_{\ell+K})$. Under the assumption that $F$ is a continuous cdf, there should be no ties in the entries of $\bf\bar{x}$} with probability one. The only other entries in $\bf z$} are the $K$ reserve prices. In practice, it is possible that there are ties in the reserve prices, in which case we add a very small noise to the reserve prices to ensure that there are no ties in the entries of $\bf z$}. Similar to $\bar{t}_{i}$, we define $\widetilde{t}_{i}=t_{\widetilde{i},\widetilde{k}}$ where $\widetilde{i}$ and $\widetilde{k}$ are such that $z_{i}=x_{\widetilde{i},\widetilde{k}}$. Further, let
@@
item is sold at or above the reserve price,
@@
Example (continued): In the example considered earlier in this subsection with $K=4$ auctions, by pooling the $4$ reserve price values with entries of $\bf\bar{x}$} and rearranging them in ascending order, we obtain
@@
Recall that the item is sold above the reserve price only in the first two auctions. By identifying the positions/ranks of $x_{m_{1},1}$, $x_{m_{2},2}$ and $x_{m_{3},3}$ (with $m_{3}=0$) in $\bf z$}, we obtain $\bar{S}=(4,9,11)$. Similarly, by identifying the positions/ranks of the entries of $\bf\bar{x}$} in $\bf z$}, we obtain ${\bf u}=(3,5,6,8,9,10,11)$.
@@
(since $F$ only has jump discontinuities at elements of $\bf z$} and is otherwise piece-wise constant). This is typical in a non-parametric setting, and we can hope/expect to only obtain estimates of the valuation distribution $F$ at the observed standing prices (including the reserve prices).
@@ 2.4 Estimation of $F$ given $\lambda$: A constraint-free reparametrization
for any constant $c>0$. To address this identifiability issue, we note that no data is observed below $z_{1}$, which is the smallest reserve price among all auctions in the data. Hence, it is quite reasonable to assume that $F(z_{1})=0$, which is equivalent to setting $\theta_{1}=1$.
@@ 3 Maximizing $Lik_{A}(\lambda,\boldsymbol{\theta})$: Coordinate ascent algorithm
where $u_{0}=0$. We now introduce notation which allows for a more compact and accessible representation of $\ln(Lik_{A})$. Recall that $\bf z$} is obtained by pooling all the $K$ reserve prices and the $\ell=\sum_{k=1}^{K}m_{k}$ ‘non-reserve’ standing prices (elements of $\bf\bar{x}$}), and $u_{j}$ represents the position of the $\bar{x}_{i}$ in $\bf z$} for $1\leq j\leq\ell+K$. In particular, $u_{\ell}$ is the position of $\bar{x}_{\ell}$, the largest ‘non-reserve’ standing price across all the $K$ auctions in $\bf z$}. In other words, $\bar{x}_{\ell}=z_{u_{\ell}}$. It is possible that $u_{\ell}<\ell+K$. For example, in settings where the reserve price in one of the auctions where the item is unsold is larger than $\bar{x}_{\ell}$, it follows that $\bar{x}_{\ell}$ is not the largest entry in $\bf z$} and $u_{\ell}<\ell+K$. With this background, we define
@@ 3.1 Coordinate-wise maximizers for $Lik_{A}(\lambda,\boldsymbol{\theta})$
This amounts to estimating $F(z_{i})$ for $i>\max(\bar{S})$ by $1$. Note that any such $z_{i}$ corresponds to a reserve price which is greater than the the largest final selling price for all auctions in the data (including auctions where the item is sold at the reserve price). Since the data offers no evidence that the support of the true valuation distribution $F$ extends up to $z_{i}$, setting the estimate of $F(z_{i})$ to $1$ indeed seems a sensible choice in this non-parametric setting.
@@ 3.2 Initial estimator of $\lambda$: Generalized method of moments
Consider first a single second price auction with reserve price $r$, and recall that $M$ denotes the number of times the selling price changes throughout the course of the auction. Our goal is to find a function $h$ such that $E[h(M)]=\lambda$. To this end, we consider the process of consumers accessing the auction whose bid value is greater than or equal to $r$. Since we are assuming that the consumers bid their true private value, it follows that the proportion of such consumers in the population of all customers is 1F(r)1-F(r), and this “thinned” process of arriving consumers with bid values greater than $r$ is a Poisson process with rate $\lambda(1-F(r))$. Let $N_{r}$ represent the total number of consumers who access the auction in the period $0,\tau]$ and have bid values greater than the reserve price $r$. Then $N_{r}\sim\mbox{Poisson}(\lambda\tau(1-F(r)))$. Moreover, given $N_{r}=n$, let $A_{i}$ ($i=1,2,\ldots,n$) represent the event that the current selling price changes after the $i^{\text{th}}$ consumer (with bid greater than $r$) accesses the auction. Let, $\textbf{1}_{A_{i}}$ be the indicator function of the occurrence of the event $A_{i}$.
@@
Here $(a)$ follows from the fact that two bids above $r$ are needed for the first change in the standing/selling price. For $(b)$, note that the arrival of $i^{\text{th}}$ consumer with bid greater than $r$ changes the selling price if and only if the corresponding bid is the highest or second highest among the $i$ reserve price exceeding bids. Note that these bid values are i.i.d. with distribution $F$ truncated above $r$. There are i(i1)i(i-1) possible choices for the joint positions of the highest and second highest bids. The $i^{th}$ bid is the highest bid for (i1)(i-1) of these choices, and the second highest bid for another (i1)(i-1) choices, leading us to $(b)$. It follows from (3.12) that
@@
If the reserve price $r$ is negligible, for example compared to the smallest final selling price seen in the data set, then it is reasonable to assume that $F(r)\approx 0$. Suppose now, we consider the data from $K$ independent second price auctions of identical copies of an item, with $M_{k}$ denoting the number of standing/selling price changes throughout the course of the $k^{th}$ auction, and with $r_{k}$ the reserve price for the $k^{th}$ auction for $1\leq k\leq K$. Let $K_{r}$ be the collection of all auction indices with negligible reserve prices. Then, it follows that from (3.13) that
@@
The details of all the steps of the coordinate ascent maximization algorithm for $Lik_{A}$ are explicitly derived above in Section 3.1. However, as mentioned previously, a crucial detail which needs to be worked out is a ‘good’ initial starting point for the algorithm. Especially for highly non-convex maximizations such as in the current setting, the choice of the initial/starting value can play a critical role in the quality of the final estimate produced by the coordinate ascent algorithm. In this section, we construct an initial estimate of $F$ based on the empirical distribution functions of both the final selling prices and the first observed bids (i.e., the price when for the first time the standing price jumps to a higher value from its respective reserve price), respectively. Note that the methodology developed in George and Hui (2012) also relies exclusively on the final selling prices. However, that approach requires the knowledge of the total number of consumers accessing each of the $K$ auctions. We do not assume this knowledge in the current setting, and need to overcome this additional challenge. Also, as stated above, we will use the first non-reserve standing prices (first observed bids) to improve the quality of our initial estimator of $F$.
@@
Step I: Construct an estimate of $F$ based on the empirical distribution function of only the final selling prices of auctions with relatively small reserve prices. First, consider a single second price auction with reserve price $X_{0}=r$. As in Section 3.2 consider the process of consumers accessing the auction whose bid value is greater than or equal to $r$, and let $N_{r}$ represents the number of such consumers that access the auction in the period $0,\tau]$. As observed in Section 3.2, this thinned process of arriving consumers is a Poisson process with rate $\lambda(1-F(r))$, and $N_{r}\sim Poisson(\lambda\tau(1-F(r)))$. We derive the conditional cdf of the final selling price $X_{M}$ given $X_{0}=r,N_{r}\geq 2$ as a function of $F(x)$. For this purpose, note that
@@
Now, coming back to our setting with $K$ independent auctions, suppose that we have multiple auctions with a given reserve price $r$ (or close to $r$) where the item is sold above the reserve price. Then based on the Glivenko-Cantelli lemma, (3.16) and (3.17), we can use the function $G_{\lambda_{r}}^{-1}$ (with an appropriate estimate of $\lambda_{r}$) applied to empirical cdf of the final selling prices of these auctions to estimate $F_{r}(x)$ for $x>r$. Setting the estimate of $F(r)$ to be zero, we can then obtain an estimate of $F(x)$ for $x>r$. Clearly, we would like to choose $r$ to be as small as possible.
With this background, let $r_{q}$ denote the $q^{th}$ quantile of the reserve prices among $\{r_{1},r_{2},\cdots,r_{K}\}$. Here $q\in(0,1)$ is a user-specified constant, and we denote the set of indices of reserve prices which lie within $0,r_{q}]$ as $V(q)$. Ideally, one would like to have a reasonable number of auctions with very small/negligible reserve prices. For example, in the Xbox data analyzed in Section 5, roughly $25\%$ of the auctions have a reserve price less than $\mathdollar 1$ (the smallest final selling price is $\mathdollar 28$). Let
@@
Step II: Incorporate the first non-reserve standing prices into the construction of the initial estimate of $F$. Consider again, to begin with, a single second price auction with reserve price $r$, and the associated thinned Poisson process of arriving consumers with valuation greater than $r$. Letting $Y_{1},Y_{2}$ represent valuations of the first two arriving consumers in the thinned process, we have
@@
be the empirical distribution function of the first non-reserve standing prices for auctions in $V(q)$. Based on (3.19), we construct the estimator $\widehat{F}_{FP}$ of $F$ as
@@
Note that $G_{FP}(z_{1})=0$, which implies that $\widehat{F}_{FP}(z_{1})=0$, where $z_{1}$ is the smallest reserve price. However, $\widehat{F}_{FP}(x)>0$ when $x$ is larger than the smallest first non-reserve standing price among auctions in $V(q)$. This smallest non-reserve standing price is often much smaller than the smallest final selling price, and hence $\widehat{F}_{FP}$ can be combined with $\widehat{F}_{SP}$ of Step I, to get a better initial estimate of $F$ as follows.
Step III: Combining the two different initial estimates, namely, $\widehat{F}_{FP}$ and $\widehat{F}_{SP}$. Let $p_{1}$ and $p_{2}$ respectively represent the largest non-reserve standing price and the smallest final selling price for auctions in $V(q)$. As discussed previously, $\widehat{F}_{SP}$ underestimates $F$ below $p_{2}$ and $\widehat{F}_{FP}$ overestimates $F$ above $p_{1}$. Let $c$ be the largest real number $\leq\min\{p_{1},p_{2}\}$ such that $\widehat{F}_{FP}(c)\leq\widehat{F}_{SP}(p_{1})$. Then, we define a function $\widehat{F}_{(0)}$ based on $\widehat{F}_{FP}$ and $\widehat{F}_{SP}$ as follows:
@@
This function $\widehat{F}_{(0)}$ in (3.21) combines the strengths of the two estimators $\widehat{F}_{FP}$ and $\widehat{F}_{SP}$, and gives a more balanced estimator of $F$ over all regions. Finally, since $G_{FP},G_{SP}$ are step functions, so are $\widehat{F}_{FP},\widehat{F}_{SP}$. It follows based on (3.21) that $\widehat{F}_{(0)}$ is a step function as well, and has jumps only at the first non-reserve standing prices and final selling prices for auctions in $V(q)$. A continuous version of this estimator, denoted by $\widehat{F}_{init}$ can be obtained by linear interpolation of the values between the jump points. Since $z_{1}\leq c$, it follows that $\widehat{F}_{(0)}(z_{1})=\widehat{F}_{FP}(z_{1})=0$.

### v3

@@ 2.2 Likelihood based on the observed data from multiple identical, non-overlapping second price auctions
Suppose we consider $K$ independent second price auctions of identical copies of an item (with possibly different starting prices $r_{1},r_{2},\ldots,r_{K}$). The observed data is $\{(M_{k},O_{k})\ ,\ \{(X_{i,k}\ ,\ T_{i-1,k})\}_{i=1}^{M_{k}}\}_{k=1}^{K}$, where $M_{k}$ denotes the number of selling price changes for the $k^{th}$ auction, $O_{k}$ denotes the indicator of the item being sold at the end of the $k^{th}$ auction, $\{X_{i,k}\}_{i=1}^{M_{k}}$ denote the selling/standing price sequence for the $k^{th}$ auction, and $\{T_{i-1,k}\}_{i=1}^{M_{k}}$ denote the sequence of intermediate waiting times between successive changes in the standing prices. Finally, let $T_{M_{k},k}=\tau-\sum_{i=0}^{M_{k}-1}T_{i,k}$, for all $k=1,2,\ldots,K$. Since the auctions are independent, it follows by Lemma 2.1 that the likelihood function of the unknown parameters $\lambda$ and $F$ for observed data values $\{(m_{k},o_{k})\ ,\ \{(x_{i,k}\ ,\ t_{i-1,k})\}_{i=1}^{m_{k}}\}_{k=1}^{K}$ is given by $Lik(\lambda,F)=\prod_{k=1}^{K}g(m_{k},o_{k},\{x_{i,k}\}_{i=1}^{m_{k}},\{t_{i,k}\}_{i=0}^{m_{k}-1})$.
@@
Note that the function $F$ is constrained to be non-decreasing. A key transformation to a constraint-free parametrization (described in Section 2.4 below) is needed to facilitate the conditional maximization of $Lik$ with respect to $F$. A crucial precursor to this re-parametrization is introduction of some new notation obtained by merging the standing prices from all the $K$ different auctions together. Let $L=\sum_{k=1}^{K}M_{k}$ denote the total number of standing/selling price changes in all the $K$ auctions. Recall that $\{(m_{k},o_{k})\ ,\ \{(x_{i,k}\ ,\ t_{i-1,k})\}_{i=1}^{m_{k}}\}_{k=1}^{K}$ denote the observed data values, and $t_{m_{k},k}=\tau-\sum_{i=0}^{m_{k}-1}t_{i,k}$ for $1\leq k\leq K$. Let $\ell$ denote the observed value of $L$. We will denote by ${\bf\bar{x}}=(\bar{x}_{1},\bar{x}_{2},\cdots,\bar{x}_{\ell})$ the arrangement/ordering of the pooled collection $\{\{x_{i,k}\}_{i=1}^{m_{k}}\}_{k=1}^{K}$ such that $\bar{x}_{1}<\bar{x}_{2}<\ldots<\bar{x}_{\ell}$; under the assumption that $F$ is a continuous cdf, there should be no ties in the entries of $\bf\bar{x}$} with probability one. In other words, we pool the standing prices from all the auctions (excluding the starting prices) and then arrange them in ascending order as $(\bar{x}_{1},\bar{x}_{2},\cdots,\bar{x}_{\ell})$. Also, for $1\leq i\leq L$, we define $\bar{t}_{i}=t_{\bar{i},\bar{k}}$ where $\bar{i}$ and $\bar{k}$ are such that $\bar{x}_{i}=x_{\bar{i},\bar{k}}$. Further, let
@@
item is sold above the starting price,
@@
Example (continued): Consider the example in Figure 2, where we have a setting with $K=4$ independent second price auctions, with starting prices $(r_{1},r_{2},r_{3},r_{4})=(10,5,13,17)$. The first auction has $m_{1}=3$ standing price changes with $(x_{1,1},\ x_{2,1},\ x_{3,1})=(12,15,19)$, the second auction has $m_{2}=4$ standing price changes with $(x_{1,2},\ x_{2,2},\ x_{3,2},\ x_{4,2})=(16,18,20,25)$, the item is sold at the starting price $r_{3}=13$ in the third auction ($m_{3}=0,o_{3}=1$), and the item is unsold in the fourth auction ($m_{4}=0,o_{4}=0$). Pooling and rearranging the standing prices (excluding starting prices) from all the auctions, we see that $\ell=3+4+0+0=7$, and
@@
Note that the auction item is sold above the starting price in the first two auctions, and the final selling prices are $x_{m_{1},1}=19$ and $x_{m_{2},2}=25$ respectively. Examining the positions of these two prices in $\bf\bar{x}$} gives us $S=\{5,7\}$. Finally, $K_{s}=\{1,2,3\}$ is the collection of auction indices where the item is sold.
@@
where $\bar{x}_{0}=0$. We now establish a final bit of notation necessary to introduce the constraint-free reparametrization of $F$. We now pool the $\ell+K$ standing prices from all the auctions (including the starting prices), i.e., $\{\{x_{i,k}\}_{i=0}^{m_{k}}\}_{k=1}^{K}$, and arrange them in ascending order as $z_{1}<z_{2}<\cdots<z_{\ell+K}$, and denote ${\bf z}:=(z_{1},z_{2},\cdots,z_{\ell+K})$. Under the assumption that $F$ is a continuous cdf, there should be no ties in the entries of $\bf\bar{x}$} with probability one. The only other entries in $\bf z$} are the $K$ starting prices. In practice, it is possible that there are ties in the starting prices, in which case we add a very small noise to the starting prices to ensure that there are no ties in the entries of $\bf z$}. Similar to $\bar{t}_{i}$, we define $\widetilde{t}_{i}=t_{\widetilde{i},\widetilde{k}}$ where $\widetilde{i}$ and $\widetilde{k}$ are such that $z_{i}=x_{\widetilde{i},\widetilde{k}}$. Further, let
@@
item is sold at or above the starting price,
@@
Example (continued): In the example considered earlier in this subsection with $K=4$ auctions, by pooling the $4$ starting price values with entries of $\bf\bar{x}$} and rearranging them in ascending order, we obtain
@@
Recall that the item is sold above the starting price only in the first two auctions. By identifying the positions/ranks of $x_{m_{1},1}$, $x_{m_{2},2}$ and $x_{m_{3},3}$ (with $m_{3}=0$) in $\bf z$}, we obtain $\bar{S}=(4,9,11)$. Similarly, by identifying the positions/ranks of the entries of $\bf\bar{x}$} in $\bf z$}, we obtain ${\bf u}=(3,5,6,8,9,10,11)$.
@@
(since $F$ only has jump discontinuities at elements of $\bf z$} and is otherwise piece-wise constant). This is typical in a non-parametric setting, and we can hope/expect to only obtain estimates of the valuation distribution $F$ at the observed standing prices (including the starting prices).
@@ 2.4 Estimation of $F$ given $\lambda$: A constraint-free reparametrization
for any constant $c>0$. To address this identifiability issue, we note that no data is observed below $z_{1}$, which is the smallest starting price among all auctions in the data. Hence, it is quite reasonable to assume that $F(z_{1})=0$, which is equivalent to setting $\theta_{1}=1$.
@@ 3 Maximizing $Lik_{A}(\lambda,\boldsymbol{\theta})$: Coordinate ascent algorithm
where $u_{0}=0$. We now introduce notation which allows for a more compact and accessible representation of $\ln(Lik_{A})$. Recall that $\bf z$} is obtained by pooling all the $K$ starting prices and the $\ell=\sum_{k=1}^{K}m_{k}$ ‘non-starting’ standing prices (elements of $\bf\bar{x}$}), and $u_{j}$ represents the position of the $\bar{x}_{i}$ in $\bf z$} for $1\leq j\leq\ell+K$. In particular, $u_{\ell}$ is the position of $\bar{x}_{\ell}$, the largest ‘non-starting’ standing price across all the $K$ auctions in $\bf z$}. In other words, $\bar{x}_{\ell}=z_{u_{\ell}}$. It is possible that $u_{\ell}<\ell+K$. For example, in settings where the starting price in one of the auctions where the item is unsold is larger than $\bar{x}_{\ell}$, it follows that $\bar{x}_{\ell}$ is not the largest entry in $\bf z$} and $u_{\ell}<\ell+K$. With this background, we define
@@ 3.1 Coordinate-wise maximizers for $Lik_{A}(\lambda,\boldsymbol{\theta})$
This amounts to estimating $F(z_{i})$ for $i>\max(\bar{S})$ by $1$. Note that any such $z_{i}$ corresponds to a starting price which is greater than the largest final selling price for all auctions in the data (including auctions where the item is sold at the starting price). Since the data offers no evidence that the support of the true valuation distribution $F$ extends up to $z_{i}$, setting the estimate of $F(z_{i})$ to $1$ indeed seems a sensible choice in this non-parametric setting.
@@ 3.2 Initial estimator of $\lambda$: Generalized method of moments
Consider first a single second price auction with starting price $r$, and recall that $M$ denotes the number of times the selling price changes throughout the course of the auction. Our goal is to find a function $h$ such that $E[h(M)]=\lambda$. To this end, we consider the process of consumers accessing the auction whose bid value is greater than or equal to $r$. Since we are assuming that the consumers bid their true private value, it follows that the proportion of such consumers in the population of all customers is 1F(r)1-F(r), and this “thinned” process of arriving consumers with bid values greater than $r$ is a Poisson process with rate $\lambda(1-F(r))$. Let $N_{r}$ represent the total number of consumers who access the auction in the period $0,\tau]$ and have bid values greater than the starting price $r$. Then $N_{r}\sim\mbox{Poisson}(\lambda\tau(1-F(r)))$. Moreover, given $N_{r}=n$, let $A_{i}$ ($i=1,2,\ldots,n$) represent the event that the current selling price changes after the $i^{\text{th}}$ consumer (with bid greater than $r$) accesses the auction. Let, $\textbf{1}_{A_{i}}$ be the indicator function of the occurrence of the event $A_{i}$.
@@
Here $(a)$ follows from the fact that two bids above $r$ are needed for the first change in the standing/selling price. For $(b)$, note that the arrival of $i^{\text{th}}$ consumer with bid greater than $r$ changes the selling price if and only if the corresponding bid is the highest or second highest among the $i$ starting price exceeding bids. Note that these bid values are i.i.d. with distribution $F$ truncated above $r$. There are i(i1)i(i-1) possible choices for the joint positions of the highest and second highest bids. The $i^{th}$ bid is the highest bid for (i1)(i-1) of these choices, and the second highest bid for another (i1)(i-1) choices, leading us to $(b)$. It follows from (3.12) that
@@
If the starting price $r$ is negligible, for example compared to the smallest final selling price seen in the data set, then it is reasonable to assume that $F(r)\approx 0$ (According to Yu et al. (2006) this is the optimal strategy for seller). Suppose now, we consider the data from $K$ independent second price auctions of identical copies of an item, with $M_{k}$ denoting the number of standing/selling price changes throughout the course of the $k^{th}$ auction, and with $r_{k}$ the starting price for the $k^{th}$ auction for $1\leq k\leq K$. Let $K_{r}$ be the collection of all auction indices with negligible starting prices. Then, it follows that from (3.13) that
@@
The details of all the steps of the coordinate ascent maximization algorithm for $Lik_{A}$ are explicitly derived above in Section 3.1. However, as mentioned previously, a crucial detail which needs to be worked out is a ‘good’ initial starting point for the algorithm. Especially for highly non-convex maximization such as in the current setting, the choice of the initial/starting value can play a critical role in the quality of the final estimate produced by the coordinate ascent algorithm. In this section, we construct an initial estimate of $F$ based on the empirical distribution functions of both the final selling prices and the first observed bids (i.e., the price when for the first time the standing price jumps to a higher value from its respective starting price), respectively. Note that the methodology developed in George and Hui (2012) also relies exclusively on the final selling prices. However, that approach requires the knowledge of the total number of consumers accessing each of the $K$ auctions. We do not assume this knowledge in the current setting, and need to overcome this additional challenge. Also, as stated above, we will use the first non-starting standing prices (first observed bids) to improve the quality of our initial estimator of $F$.
@@
Step I: Construct an estimate of $F$ based on the empirical distribution function of only the final selling prices of auctions with relatively small starting prices. First, consider a single second price auction with starting price $X_{0}=r$. As in Section 3.2 consider the process of consumers accessing the auction whose bid value is greater than or equal to $r$, and let $N_{r}$ represents the number of such consumers that access the auction in the period $0,\tau]$. As observed in Section 3.2, this thinned process of arriving consumers is a Poisson process with rate $\lambda(1-F(r))$, and $N_{r}\sim Poisson(\lambda\tau(1-F(r)))$. We derive the conditional cdf of the final selling price $X_{M}$ given $X_{0}=r,N_{r}\geq 2$ as a function of $F(x)$. For this purpose, note that
@@
Now, coming back to our setting with $K$ independent auctions, suppose that we have multiple auctions with a given starting price $r$ (or close to $r$) where the item is sold above the starting price. Then based on the Glivenko-Cantelli lemma, (3.16) and (3.17), we can use the function $G_{\lambda_{r}}^{-1}$ (with an appropriate estimate of $\lambda_{r}$) applied to empirical cdf of the final selling prices of these auctions to estimate $F_{r}(x)$ for $x>r$. Setting the estimate of $F(r)$ to be zero, we can then obtain an estimate of $F(x)$ for $x>r$. Clearly, we would like to choose $r$ to be as small as possible.
With this background, let $r_{q}$ denote the $q^{th}$ quantile of the starting prices among $\{r_{1},r_{2},\cdots,r_{K}\}$. Here $q\in(0,1)$ is a user-specified constant, and we denote by $V(q)$ the set of indices of starting prices which lie within $0,r_{q}]$. Ideally, one would like to have a reasonable number of auctions with very small/negligible starting prices. For example, in the Xbox data analyzed in Section 5, roughly $25\%$ of the auctions have a starting price less than $\mathdollar 1$ (the smallest final selling price is $\mathdollar 28$). Let
@@
Step II: Incorporate the first non-starting standing prices into the construction of the initial estimate of $F$. Consider again, to begin with, a single second price auction with starting price $r$, and the associated thinned Poisson process of arriving consumers with valuation greater than $r$. Letting $Y_{1},Y_{2}$ represent valuations of the first two arriving consumers in the thinned process, we have
@@
be the empirical distribution function of the first non-starting standing prices for auctions in $V(q)$. Based on (3.19), we construct the estimator $\widehat{F}_{FP}$ of $F$ as
@@
Note that $G_{FP}(z_{1})=0$, which implies that $\widehat{F}_{FP}(z_{1})=0$, where $z_{1}$ is the smallest starting price. However, $\widehat{F}_{FP}(x)>0$ when $x$ is larger than the smallest first non-starting standing price among auctions in $V(q)$. This smallest non-starting standing price is often considerably smaller than the smallest final selling price, and hence $\widehat{F}_{FP}$ can be combined with $\widehat{F}_{SP}$ of Step I, to get a better initial estimate of $F$ as follows.
Step III: Combining the two different initial estimates, namely, $\widehat{F}_{FP}$ and $\widehat{F}_{SP}$. Let $p_{1}$ and $p_{2}$ respectively represent the largest non-starting standing price and the smallest final selling price for auctions in $V(q)$. As discussed previously, $\widehat{F}_{SP}$ underestimates $F$ below $p_{2}$ and $\widehat{F}_{FP}$ overestimates $F$ above $p_{1}$. Let $c$ be the largest real number $\leq\min\{p_{1},p_{2}\}$ such that $\widehat{F}_{FP}(c)\leq\widehat{F}_{SP}(p_{1})$. Then, we define a function $\widehat{F}_{(0)}$ based on $\widehat{F}_{FP}$ and $\widehat{F}_{SP}$ as follows:
@@
This function $\widehat{F}_{(0)}$ in (3.21) combines the strengths of the two estimators $\widehat{F}_{FP}$ and $\widehat{F}_{SP}$, and gives a more balanced estimator of $F$ over all regions. Finally, since $G_{FP},G_{SP}$ are step functions, so are $\widehat{F}_{FP},\widehat{F}_{SP}$. It follows based on (3.21) that $\widehat{F}_{(0)}$ is a step function as well, and has jumps only at the first non-starting standing prices and final selling prices for auctions in $V(q)$. A continuous version of this estimator, denoted by $\widehat{F}_{init}$ can be obtained by linear interpolation of the values between the jump points. Since $z_{1}\leq c$, it follows that $\widehat{F}_{(0)}(z_{1})=\widehat{F}_{FP}(z_{1})=0$.